Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 61     Entry time: 01 Mar 2023, 23:25
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 1-352 pulses, 0.8 mJ, stable delivery

Optical laser delivery

  • 800 nm, 101 pulses, 282 kHz, stable delivery
  • ns laser for illumination

Achievements/Observations

  • Nanofocusing setup alignment was finished
  • Ice detection was tested

Issues

  • No major issue
ELOG V3.1.4-7c3fd00